OpenStudy (anonymous):

FINNALY i got one right....but is this onee right?? 7+12b=18b-17 -12b -12b 7+b=6b-17 -6 -6b-17 b=17??

OpenStudy (anonymous):

7+12b = 18b-17 subtract 12 b 7 = 6b-17 add 17 24 = 6b divided by 6 4 = b

OpenStudy (anonymous):

after -12b, all ur b on the left side should be gone. always subtitute ur answer back in the original question to check if it's correct.